Greenhab Report – May 24th

Dana Levin


Environmental control: Ambient

Shade cloth: on
Average temperatures: —
Low: 20C
High: 50C

Hours of supplemental light: none

Changes to crops: Per Shannon, all crops had died prior to previous crews arrival. These crops remain dead

Daily water usage for crops: n/a

Time(s) of watering for crops: n/a

Moringa research observations: n/a

Changes to research plants: n/a

Aquaponics: n/a

Narrative: Greenhab was filmed by the film crew today but no other activities took place

Support/supplies needed: n/a

Journalist Report – May 24th


Today the crew awoke to the typical breakfast and morning briefings of the day. A PR crew joined us from a neighboring habitat to film some promotional material for mission control. The crew successfully set out on their EVA and the film crew recorded aspects of habitat life and EVA operations. Unfortunately disaster befell the crew of MDRS 195 yet again. The EVA had to be aborted after an emergency call was received. The crew expertly managed the crisis and recovered all crew members and vehicles before true disaster fell. We were fortunate to have the assistance of our PR crew to help manage the disaster. But all crew remain alive so we are looking forward to tomorrow’s mission, and we remain optimistic that perhaps tomorrow our crew will escape unscathed….

Copyright © The Mars Society. All rights reserved. | Main Site